# SiniticMTError: A Machine Translation Dataset with Error Annotations for Sinitic Languages

### v1

@@ 4 Preliminary Dataset Statistics
(a) Mandarin (2009 sentences)
@@
Mistranslation 2,471 36.3% 2,481
Typography 1,723 25.6% 1,730
Omission 1,355 19.9% 1,360
Addition 470 6.9% 472
Grammar 348 5.1% 349
Spelling 244 3.6% 245
Untranslated 180 2.6% 181
Unintelligible 1 0.0% 1
(b) Cantonese (996 sentences)
Table 1: Error Category counts in Mandarin and Cantonese mt outputs, sorted by frequency.
@@
Minor 2,127 5,863
Major 1,639 949
Table 2: Error severity counts in Mandarin and Cantonese mt outputs. Minor errors are more frequent in both languages, though major errors remain substantial.
In Mandarin annotations, mistranslation, omission, and grammar errors are the most frequent error types. For Cantonese, the most common error types are omission, typography, and mistranslation. The high frequency of omission and mistranslation error types in both Mandarin and Cantonese, compared to surface level errors like spelling, untranslated, typography, and grammar, reinforces that although autoregressive language models are capable of generating well-formed and plausible sentences, they often struggle with incomplete representations of semantic and syntactic structure, (Min et al., 2025), with high sensitivity to surface form (Berglund et al., 2024; Kitouni et al., 2024) and poor semantic generalization to rare constructions (Scivetti et al., 2025). Cantonese outputs also record a much higher number of error spans per sentence compared to Mandarin, which indicates that the model performs worse in Cantonese. This likely represents the limited availability of high-quality Cantonese resources (Xiang et al., 2024). Finally, an analysis of the number of major versus minor errors shows that many errors are minor semantic or felicity issues rather than complete misinterpretations, attesting to the powerful multilingual adaptability of transformer-based language models (e.g. Liu et al., 2020; Zhu et al., 2024). Overall, these early results show both shared and language-specific challenges in Sinitic machine translation.
5 Annotation Insights
Our span-level annotations reveal insights with implications in both linguistics and natural language processing. We discuss challenges in MT systems that were also outlined in concurrent work.
@@ 5 Annotation Insights
Beyond lexical differences, Cantonese and Mandarin differ in several ways, e.g. word order and function word inventory (Zhang, 1998). The suite of aspect and feature markers and sentence-final particles differ, with Cantonese boasting a richer inventory that encodes more fine-grained nuance in tense, speaker stance or attitude (Yap and Chor, 2011; Lee, 2019). Cantonese also allows serial verb constructions (e.g. go buy eat rice as a sequence) to a greater extent than in Mandarin (Matthews, 2006). Such differences may be reflected in Table 5, where the grammar error type is much more frequent per sentence in Cantonese than in Mandarin. The difference in functional word inventory is likely a major source of such error; we observe erroneous particle uses in Cantonese mt sentences, many of which are valid particles in Mandarin yet ungrammatical in Cantonese. Examples include erroneous use of 才 in place of Cantonese particle 先 only after, Mandarin possessive 的 in place of Cantonese 嘅; and Mandarin 在 in place of Cantonese copula 喺 be at.
@@
As discussed in Section 1, MT in Cantonese and Wu Chinese face unique difficulties as a primarily a spoken language. Written Cantonese most often appears in informal contexts like texting, where conventions are inconsistent. As a result, both MT systems and Cantonese annotators face the challenge of the lack of an accepted written norm. There are several instances in the annotations where multiple written forms correspond to the same spoken word, such as 噉樣, 咁樣 like this, this way. Annotators accept both variants: 噉樣 is considered the standard form, while 咁樣 is more widely used in practice.
@@
We also observe language confusion; some machine translations in Wu or Cantonese unexpectedly output material from other languages. While it may have been unsurprising to observe language confusion within the Sinitic family, we observe one instance where flu was erroneously translated into Japanese インフルエンザ, a Katakana adoption of English word influenza. This behavior may resemble code mixing in humans, a behavior noted across bilingual infants (Lanza, 1997) and adults (Muysken, 2000). However, we believe this is closer to language confusion or interference often observed in multilingual NLP systems (Wang et al., 2020; Yu et al., 2024; Lee et al., 2025), which has been attributed to high temperature, language mismatch between representation learning and preference tuning, and under-training (Marchisio et al., 2024).
@@
6 Conclusion
@@
In this paper, we introduce SiniticMTError, a parallel English-to-Sinitic machine translation error dataset in Mandarin, Cantonese, and Wu Chinese that comprise of erroneous machine translation text, erroneous spans, their error type and severity, and a gold-label translation. The dataset is a work-in-progress; the Mandarin split is complete with 2,000 sentences; Cantonese in progress with 1,000, and Wu in its pilot stages.
@@
SiniticMTError will be one of the first human-annotated span-level resources for Wu Chinese, along with the Shanghainese Universal Dependencies dataset (ShUD; Yang, 2025) and serve as a significant addition to a small collection of Cantonese human-annotated resources. Beyond the dataset’s immediate role in error analysis and MT, we anticipate that the parallel dataset can support a wide range of downstream applications. The dataset may be adapted for several natural language understanding tasks, including language detection and linguistic acceptability judgment. Moreover, the parallel nature of the dataset makes it a promising resource for transfer learning, which allows models trained on high-resource languages to be adapted more effectively to low-resource Sinitic varieties.
@@
While we present the annotation guidelines and workflow, the annotation is incomplete at the time of submission. We plan to complete the annotation by the time the paper is to be presented, but the reported analysis is preliminary and may not be representative of our completed work.
@@
Finally, our work only spans Cantonese and Wu Chinese among over a dozen of Sinitic languages (Tang and van Heuven, 2007; Chappell, 2015), each with varying numbers of native and bilingual speakers (Norman, 1988; Eberhard et al., 2023). Future work may bootstrap our annotation guidelines and tools to expand to other Sinitic languages as well.
In the future, we plan to finish the annotation of about 2000 sentences for both Mandarin and Cantonese, as well as around 1000 sentences for Wu Chinese. We also intend to conduct experiments that showcase the utility of the datset for machine translation evaluation and error analysis. We further plan to release both the dataset and our experimental results to encourage broader adoption and future research in this area.
Acknowledgment
We thank the members of the Lee Language Lab for their valuable comments and feedback. Any errors remain our own.
Responsible Research Statement
@@ Acknowledgment
Our work introduces a dataset built from a publicly available, non-sensitive source dataset FLORES+ (Goyal et al., 2022; NLLB Team et al., 2024; Yu et al., 2024). The annotators are native speakers of the target language (Cantonese or Wu Chinese), knowledgeable about their work and the dataset’s downstream use. We internally review the work to ensure it does not contain any personally identifiable information.

### v2

@@ 3.4.  Dataset Statistics
Total 3,766 100% 1,875
Table 4: Mandarin (2009 sentences)
@@
Mistranslation 2,992 33.8% 2,134
Typography 2,329 26.3% 1,661
Omission 1,914 21.6% 1,365
Addition 600 6.8% 428
Grammar 450 5.1% 321
Spelling 326 3.7% 233
Untranslated 242 2.7% 173
Unintelligible 2 0.0% 1
Total 8,855 100% 6,316
Table 5: Cantonese (1402 sentences)
Type Count Proportion Freq / 1k
Mistranslation 46 37.4% 676
Register 42 34.1% 618
Omission 25 20.3% 368
Addition 4 3.3% 59
Spelling 3 2.4% 44
Grammar 2 1.6% 29
Typography 1 0.8% 15
Total 123 100% 1,809
Table 6: Wu Chinese (68 sentences)
Type Count Proportion Freq / 1k
Mistranslation 107 30.5% 695
Register 38 10.8% 247
Omission 29 8.3% 188
Addition 26 7.4% 169
Grammar 14 4.0% 91
Untranslated 14 4.0% 91
Other 123 35.0% 799
Total 351 100% 2,279
Table 7: Hokkien (154 sentences)
@@
Minor 2,127 7,346
Major 1,639 1,509
Table 8: Error severity counts in Mandarin and Cantonese mt outputs. Minor errors are more frequent in all languages, though major errors remain substantial.
In Mandarin annotations, mistranslation, omission, and grammar errors are the most frequent error types. For Cantonese, the most common error types are mistranslation, typography, and omission. The high frequency of omission and mistranslation error types in both Mandarin and Cantonese, compared to surface level errors like spelling, untranslated, typography, and grammar, reinforces that although autoregressive language models are capable of generating well-formed and plausible sentences, they often struggle with incomplete representations of semantic and syntactic structure, with high sensitivity to surface form (Berglund et al., 2024; Kitouni et al., 2024) and poor semantic generalization to rare constructions (Scivetti et al., 2025). Cantonese outputs also record a much higher number of error spans per sentence compared to Mandarin, which indicates that the model performs worse in Cantonese. This likely represents the limited availability of high-quality Cantonese resources (Xiang et al., 2024). Finally, an analysis of the number of major versus minor errors shows that many errors are minor semantic or felicity issues rather than complete misinterpretations, attesting to the powerful multilingual adaptability of transformer-based language models (e.g. Liu et al., 2020; Zhu et al., 2024). Overall, these results show both shared and language-specific challenges in Sinitic machine translation.
We emphasize that these observations reflect trends within the specific MT outputs used to construct SiniticMTError. Since each translation direction is represented by a single model, the reported distributions and comparisons should be interpreted as dataset-level analyses rather than general claims about English–Chinese machine translation performance. Different models may exhibit different error profiles, and future work incorporating multiple systems would allow for more system-level generalization.
4.  Annotation Insights
Our span-level annotations reveal insights with implications in both linguistics and natural language processing. Due to the small size of Wu and Hokkien pilots, we discuss insights from Mandarin and Cantonese annotations. We discuss challenges in MT systems that were also outlined in concurrent work.
@@ 4.  Annotation Insights
Beyond lexical differences, Cantonese and Mandarin differ in several ways, e.g. word order and function word inventory (Zhang, 1998). The suite of aspect and feature markers and sentence-final particles differ, with Cantonese boasting a richer inventory that encodes more fine-grained nuance in tense, speaker stance or attitude (Chor, 2018; Lee, 2019). Cantonese also allows serial verb constructions (e.g. go buy eat rice as a sequence) to a greater extent than in Mandarin (Matthews, 2006). Such differences may be reflected in Table 4 and Table 5, where the grammar error type is much more frequent per sentence in Cantonese than in Mandarin. The difference in functional word inventory is likely a major source of such error; we observe erroneous particle uses in Cantonese mt sentences, many of which are valid particles in Mandarin yet ungrammatical in Cantonese. Examples include erroneous use of 才 cai2 in place of Cantonese particle 先 sin1 only after, Mandarin possessive 的 de in place of Cantonese 嘅 ge3; and Mandarin 在 zai4 in place of Cantonese copula 喺 hai2 be at.
@@
As discussed in Section 1, machine translation in Cantonese and Wu Chinese face unique difficulties as a primarily a spoken language with only a short history of writing (Snow, 2004, 2008). Written Cantonese most often appears in informal contexts like texting, where conventions are inconsistent. As a result, both MT systems and Cantonese annotators face the challenge of the lack of an accepted written norm. There are several instances in the annotations where multiple written forms correspond to the same spoken word, such as 噉樣 gam2joeng2, 咁樣 gam2joeng2 like this, this way. Annotators accept both variants: 噉樣 is considered the standard form, while 咁樣 is more widely used in practice.
However, such lack of standardization is not to say that there is no systematicity in Cantonese or Wu orthography and by extension, machine translation error annotation. As noted in previous work (Lu et al., 2024b), orthographies can reliably be converted from one to another; each error we annotate is similarly annotatable across orthography. In SiniticMTError, we annotate using Hanzi and Hanzi-derived systems and their general conventions with which annotators are most familiar. We do not annotate acceptable alternative orthography (cf. color and colour in English).
@@
We also observe language confusion; some machine translations in Wu or Cantonese unexpectedly output material from other languages. While it may have been unsurprising to observe language confusion within the Sinitic family (e.g. Cantonese 下 haa6 for 落 lok6 down, following Mandarin 下 xia4 down), we observe one instance where flu was erroneously translated into Japanese インフルエンザ infuruenza, a Katakana adoption of Italian-derived-English word influenza. This behavior superficially resembles code mixing observed in bilingual speakers (Lanza, 1997; Muysken, 2000). However, in multilingual NLP systems it is more commonly interpreted as language confusion or interference (Wang et al., 2020; Yu et al., 2024; Lee et al., 2025), which has been attributed to high temperature, language mismatch between representation learning and preference tuning, and under-training (Marchisio et al., 2024).
@@
5.  Language Model Baselines
In addition to dataset statistics, we report language model baselines on SiniticMTError. To benchmark the effectiveness of different LLMs in detecting translation errors, we adopt two evaluation schemes inspired by MQM-APE Lu et al. (2025).
5.1.  Error Span Evaluation
Following MQM-APE, we evaluate the span-level agreement between LLM-generated error annotations and human references using Span Precision (SP) and Major Precision (MP). More formally, for any single error span $e$, we define $P(e)=\{i,i+1,...,j\}$, where $i$ is the start of the marked error span and $j$ is the end of the error span. For a collection of error spans $E=\{e_{1},...,e_{n}\}$, we also define $P(E)=\bigcup^{n}_{j=1}P(e_{j})$. Then, SP and MP are defined as follows:
SP $=\frac{P(E)\cap P(\hat{E})}{P(\hat{E})}$ (1)
MP $=\frac{P(E_{maj})\cap P(\hat{E}_{maj})}{P(\hat{E}_{maj})}$ (2)
where $E$ is a collection of gold error spans, $\hat{E}$ is a collection of LLM generated error spans, and $E_{maj}$ denotes the subset of errors that are identified as major errors.
SP measures precision across all error spans predicted by the model, whereas MP focuses exclusively on major error spans, which have the greatest impact on translation quality under the Multidimensional Quality Metrics (MQM) Lommel et al. (2013) framework. This metric measures the quality of error span alignment, as it verifies if the indices of the identified error spans align with the gold error spans’ indices.
Independently of the span index matching, we also compare LLM and human-derived segment-level MQM scores using correlation metrics.
We convert annotations to numeric MQM via the standard weighting scheme (e.g., Minor=1, Major=5) (Freitag et al., 2021) using the official converter from mt-metrics-eval [^6] https://github.com/google-research/mt-metrics-eval, consistent to previous works (Lu et al., 2024c; Kocmi and Federmann, 2023; Fernandes et al., 2023).
We then compute segment-level alignment between human and LLM-derived MQM scores via Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$, which are standard correlation-based metrics widely used in QE and MT evaluation research (Freitag et al., 2021; Stefanik et al., 2021).
Model en-zh en-yue
SP MP SP MP
GPT-4o 40.5 26.8 51.7 14.9
Gemini-2.5-pro 40.9 25.9 59.7 19.8
Gemma-3-12B-it 28.1 25.7 41.4 18.9
Microsoft/Phi-4 29.3 22.8 48.2 12.2
Qwen3-14B 44.3 29.7 57.8 14.0
Table 9: Span Precision (SP) and Major Precision (MP) across en-zh and en-yue. Bold indicates best performance under a metric.
Model Task $r$ $\rho$ $\tau$
GPT-4o en-zh 0.43 0.43 0.33
en-yue 0.19 0.21 0.16
Gemini-2.5-pro en-zh 0.45 0.46 0.36
en-yue 0.35 0.32 0.24
Gemma-3-12B-it en-zh 0.34 0.34 0.27
en-yue 0.07 0.07 0.05
Microsoft/Phi-4 en-zh 0.29 0.27 0.21
en-yue 0.00 0.02 0.02
Qwen3-14B en-zh 0.40 0.41 0.32
en-yue 0.15 0.09 0.01
Table 10: Segment-level MQM score correlation coefficients (Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$) across two language pairs. Bold indicates best performance under a metric. Language codes en, zh, and yue correspond to English, Mandarin, and Cantonese, respectively.
5.2.  Index Correctness within LLMs
To ensure that the indices of LLM-predicted error spans are consistent with the character positions in the mt string, we introduce a lightweight tag-to-span generation protocol (tag2span). The model outputs a tagged version of the machine translation together with a structured error list, following a simple JSON schema. Each error region in the translation is enclosed by a unique identifier tag such as <spanN> ... </spanN>. We then deterministically parse <spanN> ... </spanN> boundaries to recover $(\textit{start\_index},\textit{end\_index})$ pairs in the original mt. This enforces an unambiguous alignment between model-declared error regions and character offsets, eliminating off-by-one and tokenization-related drift during post-processing.
We also employ the zero-denominator policy to ensure correctness. Since SP and MP are precision measures, denominators can be zero. If $\left|P(\hat{E})\right|=0$, then SP is undefined, similar for MP. In such cases we report the score as N/A. We do not impute 0 for undefined precision to avoid conflating “predicts nothing” with “predicts many positions but entirely wrong.”
5.3.  Interpretation
As shown in Tables 9 and 10, across all evaluated models on the en-zh and en-yue setting, both span-level and correlation-based metrics remain modest. SP and MP values range mostly between 0.25–0.45, while segment-level correlations ($r$, $\rho$, $\tau$) between LLM-derived and human MQM scores stay below 0.5. These consistently limited scores indicate that current open-source LLMs, such as Gemma-3-12B-it, Phi-4, and Qwen3-14B, lack the specialized training required for fine-grained translation error analysis. Furthermore, performance on closed-source LLMs seems limited as well, as both segment-level analysis and MQM score correlation scores are in similar ranges across language pairs.
In addition to this domain-specific limitation, another contributing factor is the low-resource nature of the benchmarked language pairs. Languages such as Cantonese are severely underrepresented in existing MT datasets and LLM pretraining corpora, leading to weaker cross-lingual representations and reduced sensitivity to fine-grained translation schemes. This scarcity of high-quality bilingual data makes it difficult for models to accurately identify context-dependent or culturally specific translation errors.
Together, these findings underscore the necessity of our newly constructed dataset. By providing high-quality human annotations for underrepresented language pairs and explicit error span annotations, our dataset fills a gap in the current multilingual translation and evaluation ecosystem.
It not only enables the systematic study of translation error patterns in low-resource settings but also provides a foundation for fine-tuning and aligning future LLMs toward more accurate, human-consistent translation quality assessment.
6.  Conclusion
@@
In this paper, we introduce SiniticMTError, a dataset of machine translation errors, comprising parallel English–Mandarin, English–Cantonese, and English–Wu Chinese splits, as well as a separate Mandarin–Hokkien component. Each entry contains an erroneous machine translation text, erroneous spans, their respective error type and severity, and a gold-label translation. The Mandarin split contains 2,009 annotated sentences; Cantonese 1,402, Wu 68, and Hokkien 154.
@@
SiniticMTError is one of the first human-annotated span-level MT error resources for Wu Chinese, complementing other emerging resources such as the Shanghainese UD dataset (Yang, 2025). It also serves as a significant addition to a small collection of Cantonese and Hokkien human-annotated resources. In addition, our baseline experiments show that even the strongest open and closed-source LLMs achieve modest span-precision and correlation scores, particularly. These results demonstrate both the difficulty of automatic error detection in low-resource Sinitic MT and the potential of SiniticMTError to serve as a training and evaluation benchmark for future models. Beyond the dataset’s immediate role in error analysis and MT, we anticipate that the parallel dataset can support a wide range of downstream applications. The dataset may be adapted for several natural language understanding tasks, including language detection and linguistic acceptability judgment (e.g. Min et al., 2025). Moreover, the parallel nature of the dataset makes it a promising resource for transfer learning, which allows models trained on high-resource languages to be adapted more effectively to low-resource Sinitic varieties.
@@
While we present datasets in four languages, two splits (Wu and Hokkien) are pilot annotations with only a small number of annotated sentences. Future work may benefit from an extension of these datasets.
@@
Finally, our work only spans Mandarin, Cantonese, Wu Chinese, and Hokkien among over a dozen of Sinitic languages (Tang and van Heuven, 2007; Chappell, 2015), each with varying numbers of native and bilingual speakers (Norman, 1988; Eberhard et al., 2023). Future work may bootstrap our annotation guidelines and tools to expand to other Sinitic languages as well.
Ethics Statement
@@
Our work introduces a dataset built from the publicly available FLORES+ (Goyal et al., 2022; NLLB Team et al., 2024; Yu et al., 2024). The annotators are native speakers of the target language (Mandarin, Cantonese or Wu Chinese), knowledgeable about their work and the dataset’s downstream use. We internally review the annotations to ensure it does not contain any sensitive or personally identifiable information.